SOLVE.
5/8x + 1/16x = 3/16 + x
(these are all fractions)
HELP, how do I do this? Do i have to change the 5/8 to 10/16x?

Answer by stanbon(75887)

(5/8)x + (1/16)x = (3/16) + x
Multiply thru by 16 to get:
10x + x = 3 + 16x
5x = -3
x = -3/5
=============
Cheers,
Stan H.